# P3H-25-058, TTP25-027

### v1

@@ 3.3 Emerging flavor mixing angles
Compared with invoking unknown dynamics above GUT scale, a more convincing approach is to add more terms to Eq (22) and change the evolution below $M_{\text{GUT}}$. This can be achieved by requiring more light scalar components in $126_{H}$, such as di-quarks Arnold:2009ay ; Giudice:2011ak ; Patel:2022nek , who also speed up the running for all Yukawa couplings. As a result, $\epsilon^{bs}$ increases more rapidly when from UV to IR and allows a much lower value at GUT scale. Interestingly, the di-quarks are also relevant for anomalies. They induce effective four-quark interactions that may account for the observed CP asymmetry in $D_{0}\rightarrow\pi^{+}\pi^{-}$ and $D_{0}\rightarrow K^{+}K^{-}$ decays LHCb:2019hro ; LHCb:2022lry ; Altmannshofer:2012ur ; Iguro:2024uuw . We leave a detailed analysis on this direction for future work.
@@
The behavior of left-handed down-type quark mixing $\epsilon^{sb}$ is similar. However, there is an additional requirement to preserve small $V_{cb}\sim V_{ts}\sim 0.04$: $b_{L}-s_{L}$ mixing should be well aligned with $t_{L}-c_{L}$ mixing. This further requires $\epsilon^{sb}=\epsilon^{ct}$ to good precision at $M_{\text{GUT}}$, which can be protected by the $SU(2)_{R}$ symmetry that connects $t_{R}H_{u}$ to $b_{R}H_{d}$. At low energies, the unbroken $SU(2)_{L}$ gauge symmetry ensures the evolution for $\epsilon^{sb}$ and $\epsilon^{ct}$ remains identical. We think it is worth offering a further comment here. Explaining $B$ anomalies requires misaligned bsb-s mixing for $H_{d}$ and LQ Yukawa couplings. In the SM, however, neither $b_{L}-s_{L}$ nor $b_{R}-s_{R}$ mixing is physical; the only observable is the difference between $b_{L}-s_{L}$ and $t_{L}-c_{L}$ mixing. In other words, the SM quark mixing pattern and beyond SM flavor transition could have intrinsically different origins, well-separated by symmetries.
@@ 4 Conclusion and outlook
In this article, we demonstrate that the TeV-scale LQs addressing the $B$ anomalies can also resolve the wrong $b-\tau$ mass relation in GUTs with simple Yukawa sectors. Among the six types of LQs contained in $126_{H}$, $\widetilde{S}_{1},\widetilde{R}_{2},S_{3}$ can account for the $b\rightarrow c,s$ anomalies while remaining compatible with grand unification. Fixing exact Yukawa unification at $10^{16}$ GeV and $m_{\widetilde{S}_{1}}=m_{\widetilde{R}_{2}}=m_{S_{3}}=1~\text{TeV}$, the running masses $m_{t},m_{b},m_{\tau}$ can match well with low-scale observations. This is a predictive result because the theory contains only two arbitrary inputs at leading-log level: $\tan\beta$ and $y_{t}(M_{\text{GUT}})$. We also show that scenarios without $\widetilde{S}_{1}$ or $\widetilde{R}_{2}$ can still lead to successful third-generation Yukawa unification. This implies that the successful prediction of mass ratios is not tied to a specific choice of LQ types. Although our minimal SO(10) theory predicts no flavor mixing, we demonstrate that the flavor conservation solution becomes unstable once the TeV-scale LQs are introduced. Nevertheless, flavor mixing angles large enough to explain the anomalies emerge only when the $126_{H}$ spectrum is further modified.
Although the central goal of this work is to illustrate $b-\tau$ unification through a simplified setup, identifying an explicit example that yields an improved global fit for $B$ anomalies would certainly strengthen our conjecture. We note, however, that such a fit introduces additional free inputs. In particular, $m_{\widetilde{S}_{1}},m_{\widetilde{R}_{2}},m_{S_{3}}$ all become relevant parameters and can no longer be simply fixed at 1 TeV, as they directly suppress the bottom quark transition amplitudes. Another new relevant set of parameters concerns to which degree the Yukawa sector at the GUT scale deviates from exact Yukawa unification, since the RG evolution of the mixing angles shows a sensitive dependence on initial conditions.
@@
At the end we remark that the physical implication of this work goes beyond unified theories. The original Froggatt-Nielsen paper Froggatt:1978nt introduced two distinct ideas to address the SM fermion mass pattern: (a) RG running effects and (b) UV model building. The idea (a) is further developed using the framework of fixed points Pendleton:1980as ; Hill:1980sq , suggesting that the IR, rather than the UV, gauge structure determines the SM flavor parameters. However, the RG equations derived from SM alone do not lead to a simple UV flavor structure, which brings more efforts to idea (b), UV model building. Thanks to the recent progress at the high-intensity frontier and in lattice QCD calculations, particularly the successful extraction of $R(D^{(*)})$, we now have data-driven motivations to extend the SM into a larger IR theory with LQs at the TeV-scale. The situation has changed compared to 20 years ago and the RG approach to the flavor puzzle deserves new attention. Although light LQs bring a new hierarchy problem, we argue that the RG approach is somehow a more promising path than UV model-building, because the emergence of complexity from simple structures at small distance scales is a universal phenomenon and not limited to particle physics Anderson:1972pca . Well-known examples can be found in condensed matter Wilson:1974mb or other complex systems Wolfram:1985qr ; Wolfram:2020jjc .
@@ Appendix A Model independence
Figure 6: RG evolution in two alternative scenarios to demostrate model independence. Top: $S_{3}+\widetilde{R}_{2}$ with $y_{t}(M_{\text{GUT}})=0.62,~\tan\beta=49$. Bottom: $S_{3}+\widetilde{S}_{1}$ with $y_{t}(M_{\text{GUT}})=0.59,~\tan\beta=46$. All others setting are same as in Figure 2.
We show scenarios with $S_{3}+\widetilde{R}_{2}$ and $S_{3}+\widetilde{S}_{1}$ in Figure 6 to demonstrate that our results do not depend on the specific choice of LQ type but are rather universal. We take the same technical setup as the three-LQ scenario, but take $y_{t}(\mu_{\text{GUT}})=0.62,~\tan\beta=49,~y_{1}\equiv 0$ for the $S_{3}+\widetilde{R}_{2}$ case, and $y_{t}(\mu_{\text{GUT}})=0.59,~\tan\beta=46,~y_{2}\equiv 0$ for the $S_{3}+\widetilde{S}_{1}$ case. Interestingly, the $S_{3}+\widetilde{R}_{2}$ scenario, which may serve as the minimal LQ model to address the B-anomaly, yields a better TeV-scale match than the three-LQ result. In contrast, $S_{3}+\widetilde{S}_{1}$ scenario might be unable to fully account for the $R(D^{(*)})$ anomaly without violating the $B\rightarrow K^{*}\overline{\nu}\nu$ constraint, its matching result is also slightly worse. Despite this, all scenarios provide consistent indication that TeV-scale LQs make the $b-\tau$ mass relationship no longer a major concern for minimal SO(10).

### v2

@@ 3.3 Emerging flavor mixing angles
Instead of invoking unknown dynamics above the GUT scale, a more convincing approach is to add more terms to Eq. (22) and change the evolution below $M_{\text{GUT}}$. This can be achieved by requiring more light scalar components in $126_{H}$, such as di-quarks Arnold:2009ay ; Giudice:2011ak ; Patel:2022nek ; Crivellin:2023saq , which also speed up the running of all Yukawa couplings. As a result, $\epsilon^{bs}$ increases more rapidly on the path from UV to IR and permits a much lower value at the GUT scale. From a phenomenological perspective, the di-quarks have also been discussed to explain the $b\rightarrow s\ell\ell$ anomalies Crivellin:2023saq and they may furthermore contribute to the interpretation of the observed CP asymmetry in $D_{0}\rightarrow\pi^{+}\pi^{-}$ and $D_{0}\rightarrow K^{+}K^{-}$ decays LHCb:2019hro ; LHCb:2022lry ; Altmannshofer:2012ur ; Iguro:2024uuw . We leave a detailed analysis in this direction for future work.
@@
The behavior of left-handed quark mixing $\epsilon^{sb}$ and $\epsilon^{ct}$ is similar. However, there is an additional element ensuring the preservation of the small $V_{cb}\sim-V_{ts}\sim 0.04$ in the RG evolution: $b_{L}-s_{L}$ mixing should be well aligned with $t_{L}-c_{L}$ mixing. This further requires $\epsilon^{sb}=\epsilon^{ct}$ to a good precision at $M_{\text{GUT}}$, which can be achieved and protected by $SU(2)_{R}$ symmetry that connects $t_{R}H_{u}$ to $b_{R}H_{d}$. Since $SU(2)_{L}$ gauge symmetry stays unbroken until low energies, the evolution for $\epsilon^{sb}$ and $\epsilon^{ct}$ is identical. We think that this feature is worth a general comment: Explaining $B$ anomalies requires misaligned bsb-s mixing for $H_{d}$ and LQ Yukawa couplings. In the SM, however, neither $b_{L}-s_{L}$ nor $b_{R}-s_{R}$ mixing is physical; the only observable is the difference between $b_{L}-s_{L}$ and $t_{L}-c_{L}$ mixing. In other words, the SM quark mixing pattern and beyond-SM flavor structure could have intrinsically different origins, with different relevant symmetries.
@@ 4 Conclusion and outlook
In this article, we demonstrate that TeV-scale LQs addressing the $B$ anomalies can also resolve the wrong $b-\tau$ mass relation in GUTs with simple Yukawa sectors. Among the six types of LQs contained in $126_{H}$, $\widetilde{S}_{1},\widetilde{R}_{2},S_{3}$ can explain the $b\rightarrow c,s$ anomalies while remaining compatible with grand unification. Fixing exact Yukawa unification at $10^{16}$ GeV and $m_{\widetilde{S}_{1}}=m_{\widetilde{R}_{2}}=m_{S_{3}}=1~\text{TeV}$, the running masses $m_{t},m_{b},m_{\tau}$ can match well with low-scale observations. This is a predictive result because the theory contains only two arbitrary inputs at leading-log level: $\tan\beta$ and $y_{t}(M_{\text{GUT}})$. We also show that scenarios without $\widetilde{S}_{1}$ or $\widetilde{R}_{2}$ can still lead to successful third-generation Yukawa unification. This implies that the successful prediction of mass ratios is not tied to a specific choice of LQ types. Although our minimal SO(10) theory predicts no flavor mixing, we demonstrate that the flavor-conserving solution becomes unstable when perturbed by small flavor-violating terms at the GUT scale (or slightly above) once TeV-scale LQs are introduced.
Our central goal is to show how Yukawa unification is improved by the LQs. Further identifying an explicit example that helps to understand the $B$ anomalies certainly strengthens our approach. A relevant set of new parameters describe the deviations from a minimal Yukawa sector is needed to obtain flavor mixing. As an important generic result, we find that flavour violation in the LQ couplings increases by RG effects while small flavour mixing in the SM Higgs Yukawa sector stays small under the RG flow by SU(2) symmetry. Nevertheless, flavor mixing angles large enough to explain the anomalies can emerge from tiny, SM-like flavor violation at the GUT scale only when the $126_{H}$ spectrum is further modified. In particular, the light colored states — such as di-quarks and color-octets — are typically required to further enhance the RG evolution of the flavor mixing angles determining the flavor-changing LQ couplings in the IR. These light colored states living in $126_{H}$ can simultaneously improve successful gauge coupling unification Goto:2023qch , because the light $S_{3}$ LQ accelerates the running of $g_{2}$, which must be balanced by a corresponding speed-up of $g_{3}$ induced by other light colored states.
The successful $b-\tau$ unification and enhancement for flavor mixings raises the question whether TeV-scale LQ can improve the unification of all three generation SM fermions. In principle, light LQs can reduce the number of Higgs multiplets needed compared to the usual GUT set-up. The TeV-scale theory then contains less free parameters compared to the generic frameworks, and some LQ couplings to the 2nd generation fermions may be related to known fermion masses and mixings. This could result in correlations among flavor observables related to lighter fermions, for instance, g2g-2 of the muon and $CP$ violation for $D$ mesons. However, second-generation fermion masses and related observables are not robustly predicted, because they can be significantly changed by tiny corrections to the minimal theory.
@@
Finally we remark that the physical implication of this work goes beyond unified theories. The original Froggatt-Nielsen paper Froggatt:1978nt introduced two distinct ideas to address the SM fermion mass pattern: (a) RG running effects and (b) UV model building. The idea (a) is further developed using the framework of fixed points Pendleton:1980as ; Hill:1980sq , suggesting that the IR, rather than the UV, structure of the theory determines the SM flavor parameters. However, the RG equations of the SM alone do not lead to a simple UV flavor structure, which underpins idea (b), UV model building. Thanks to the recent progress at the high-intensity frontier by LHCb and $B$ factories and in lattice QCD calculations, particularly the successful measurements and predictions of $R(D^{(*)})$, we now have data-driven motivations to extend the SM to a larger IR theory with LQs with TeV-scale masses. The situation has changed compared to 20 years ago and the RG approach to the flavor puzzle deserves new attention. Although light LQs bring a new hierarchy problem, we argue that the RG approach is somehow a more promising path than UV model-building, because the emergence of complexity from simple structures at small distance scales, as we observe in our RG evolution, is a universal phenomenon and not limited to particle physics.Well-known examples can be found in condensed matter Wilson:1974mb or other complex systems Wolfram:1985qr ; Wolfram:2020jjc .
@@ Appendix A Model independence
Figure 6: RG evolution in two alternative scenarios to demonstrate model independence. Top: $S_{3}+\widetilde{R}_{2}$ with $y_{t}(M_{\text{GUT}})=0.62,~\tan\beta=49$. Bottom: $S_{3}+\widetilde{S}_{1}$ with $y_{t}(M_{\text{GUT}})=0.59,~\tan\beta=46$. All others setting are same as in Figure 2.
We show scenarios with $S_{3}+\widetilde{R}_{2}$ and $S_{3}+\widetilde{S}_{1}$ in Figure 6 to demonstrate that our results do not depend on the specific choice of LQ type but are rather universal. We take the same technical setup as the three-LQ scenario, but now find successful Yukawa unification for $y_{t}(\mu_{\text{GUT}})=0.62,~\tan\beta=49,~y_{1}\equiv 0$ in the $S_{3}+\widetilde{R}_{2}$ case, and for $y_{t}(\mu_{\text{GUT}})=0.59,~\tan\beta=46,~y_{2}\equiv 0$ in the $S_{3}+\widetilde{S}_{1}$ case. Interestingly, the $S_{3}+\widetilde{R}_{2}$ scenario, which may serve as the minimal LQ model to address the B anomaly, yields a better TeV-scale match than the three-LQ result. In contrast, $S_{3}+\widetilde{S}_{1}$ scenario might be unable to fully account for the $R(D^{(*)})$ anomaly without violating the $B\rightarrow K^{*}\overline{\nu}\nu$ constraint, its matching result is also slightly worse. Despite this, all scenarios provide consistent indication that TeV-scale LQs make the $b-\tau$ mass relationship no longer a major concern for minimal SO(10).